A mineshaft has an elevator hung from a single steel-wire cable of diameter 2.5 cm. Young's modulus of the cable is 10×1010N/m2. When the cable is fully extended, the end of the cable is 900 m below the support.

How much does the fully extended cable stretch when 1200 kg of ore is loaded into the elevator?
Answer #1

d =2.5cm , r = 1.25 cm , Y = 10*10^10 N/m^2, m =1200 kg

Young modulus Y = F/A/DL/L = mg*L/A*DL

10*10^10 = (1200*9.8*900)/(3.14*1.25*1.25*10^-4*DL)

DL = 0.216 m